# ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM

### v1

@@ 4.1 Text-centric Head Selection
1:Image $\mathbf{I}$, Text-centric heads $\mathcal{H}_{\textsc{pos}}$ (from Algorithm 1), Critical vis-token count $\kappa_{\textsc{vis}}$, Steer strengths $\alpha_{\textsc{pos}},\alpha_{\textsc{neg}}$, Contrastive weight $\alpha$, Truncation threshold $\beta$, MLLM with $L$ layers and $H$ heads per layer
2:$p_{\theta}^{\text{final}}$ (final logits from ASCD)
3:Step 1: Forward Pass with Positive Steering
4:for $l\leftarrow 1$ to $L$ do
5:     for $h\leftarrow 1$ to $H$ do
6:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
7:         if $(l,h)\in\mathcal{H}_{\textsc{pos}}$ then
8:              $\mathbf{A}^{(l)}_{h}\;\leftarrow\;\mathbf{A}^{(l)}_{h}\;+\;\alpha_{\textsc{pos}}\,|\mathbf{A}^{(l)}_{h}|$
9:         end if
10:     end for
11:     Normalize $\mathbf{A}^{(l)}$ and continue
12:end for
13:Obtain logits $p_{\theta}^{\text{pos-steered}}$
14:Step 2: Forward Pass with Negative Steering
15:for $l\leftarrow 1$ to $L$ do
16:     for $h\leftarrow 1$ to $H$ do
17:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
@@
18:         Identify top-$\kappa_{\textsc{vis}}$ critical visual tokens $\mathcal{V}$
19:         for all $v\in\mathcal{V}$ do
20:              $\mathbf{A}^{(l)}_{h}(v)\;\leftarrow\;\mathbf{A}^{(l)}_{h}(v)\;-\;\alpha_{\textsc{neg}}\,|\mathbf{A}^{(l)}_{h}(v)|$
21:         end for
22:     end for
23:     Normalize $\mathbf{A}^{(l)}$ and continue
24:end for
25:Obtain logits $p_{\theta}^{\text{neg-steered}}$
26:Step 3: Contrastive Decoding with Truncation
27:$p_{\theta}^{\text{raw}}\;\leftarrow\;(1+\alpha)\,p_{\theta}^{\text{pos-steered}}\;-\;\alpha\,p_{\theta}^{\text{neg-steered}}$
[…]
@@
29:$p_{\theta}^{\text{final}}\;\leftarrow\;p_{\theta}^{\text{raw}}.\text{masked\_fill}(p_{\theta}^{\text{pos-steered}}<\text{cutoff},-\infty)$
30:return $p_{\theta}^{\text{final}}$
4.2 Attention Steering
@@
Text-centric Head Awareness and Critical Visual Token Selection. As shown in Figure 4, we refine our method by incorporating text-centric head selection for positive steering and critical token identification for negative steering. Specifically, given the selected text-centric heads, we positively steer them by increasing their attention weights with a strength of $\alpha_{\textsc{pos}}$. Figure 5a highlights how targeted steering in text-centric heads improves the positive steering effectiveness. Simultaneously, we apply negative steering to the top $\kappa_{\textsc{vis}}$ most critical visual tokens – those receiving the highest aggregate attention across heads – reducing their attention scores by $\alpha_{\textsc{neg}}$ across all heads. Through this strategy, we deliberately obscure only the most pivotal cues – this targeted suppression is sufficient to induce a strong hallucination effect in the negative branch, leading to improved contrastive decoding compared to a blanket suppression of all visual tokens. In Figure 5b, we demonstrate the impact of selectively applying negative steering to critical visual tokens.
@@
Figure 5: Comparative effectiveness of selective attention steering. (a): Positive steering applied only to text-centric heads (32 heads with the highest text-to-visual ratio) outperforms random or blanket head selection across various decoding strategies (Greedy, Nucleus, Beam), leading to higher POPE Accuracy and F1. (b): Negative steering focused on a small subset of critical visual tokens, integrated with contrastive decoding, significantly reduces CHAIR metrics (less hallucination) and boosts POPE metrics compared to randomly suppressing visual tokens of the same number. These results validate that targeted attention modulation on text-centric heads (for positive steering) and critical visual tokens (for negative steering) yields stronger hallucination mitigation and more grounded responses.
Integration with Contrastive Decoding with Truncation. Building on the attention-steering process, we first obtain two output distributions: $p_{\theta}^{\text{pos-steered}}$ from the positively steered branch and $p_{\theta}^{\text{neg-steered}}$ from the negatively steered branch. We then combine these into contrastive decoding with a truncation mechanism, as detailed in the Step 3 of Algorithm 2. This process not only reinforces visually grounded predictions but also effectively mitigates the influence of spurious textual biases.
5 Experiments
@@
To evaluate the effectiveness of our attention-steerable contrastive decoding framework in mitigating hallucinations in MLLMs, we conduct a range of experiments. This includes three diverse benchmarks – CHAIR, POPE, and MMHal-Bench – each designed to assess different aspects of object hallucinations. To ensure the broad applicability and robustness of our approach, we also test it on three representative models – LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, and employ three different decoding strategies: greedy search, nucleus sampling, and beam search. Details of the experimental settings are provided in Appendix B. Furthermore, we evaluate performance on standard VQA benchmarks including MMMU, MM-VET, ScienceQA, TextVQA, and GQA to verify that the proposed method preserves – rather than diminishes – the model’s original visual understanding.
@@
Model Decoding Method CHAIRs (↓) CHAIRi (↓)
LLaVA-1.5 7B greedy Orig 53.2 13.5
VCD 56.8 15.2
ICD 52.8 13.2
ASCD 35.6 8.6
nucleus Orig 59.0 17.4
VCD 59.8 16.6
ICD 57.4 15.6
ASCD 43.6 11.3
beam Orig 54.8 15.3
VCD 58.8 16.4
ICD 52.6 13.9
ASCD 40.8 10.1
LLaVA-NeXT 7B greedy Orig 31.6 7.5
VCD 37.2 9.7
ICD 32.8 8.4
ASCD 21.8 7.0
nucleus Orig 30.4 8.0
VCD 40.4 10.4
ICD 39.4 9.9
ASCD 21.2 6.7
beam Orig 34.0 8.5
VCD 36.6 9.1
ICD 31.8 7.6
ASCD 21.0 6.5
Phi2-SigLIP greedy Orig 29.0 6.9
VCD 39.4 9.6
ICD 33.4 7.7
ASCD 21.8 5.4
nucleus Orig 36.0 9.8
VCD 36.0 8.1
ICD 37.0 9.4
ASCD 26.0 8.0
beam Orig 30.4 6.9
VCD 36.0 8.4
ICD 31.0 7.0
ASCD 24.6 5.7
Table 1: CHAIR Evaluation Results. Lower CHAIRs and CHAIRi values indicate better performance in reducing hallucination. The best values for each metric within a model-decoding combination are highlighted in bold.
@@
It is important to note that current benchmarks for evaluating multimodal models are highly variable. For example, baseline models such as LLaVA-1.5 7B often report different metric values between different papers. Moreover, the CHAIR metric relies on random image sampling, which further complicates direct comparisons between papers. To address these issues, we faithfully reproduced both VCD and ICD using the parameters specified in their original papers and repositories, ensuring that our evaluations are conducted under consistent conditions. This allows for a more reliable comparison between our method and existing approaches.
Model Decoding Method Popular Acc Popular F1 Random Acc Random F1 Adv. Acc Adv. F1 Avg. Acc Avg. F1
@@
CHAIR. Table 1 shows the CHAIR metrics (CHAIRs and CHAIRi), which measure object hallucination in image captioning. Across all models and decoding strategies, ASCD consistently achieves lower CHAIR values than Orig, VCD, or ICD, which illustrates ASCD’s effectiveness at mitigating object-level hallucinations.
POPE. Table 2 reports the accuracy and F1 scores under the POPE evaluation, which probes object presence with random, popular, and adversarial queries. Higher values indicate fewer hallucinations. Again, ASCD achieves the best or near-best performance in all cases. These gains persist across different model architectures, suggesting that attention steering is robust to model size and design variations.
MMHal-Bench. Figure 7 illustrates the radar charts of MMHal-Bench results for LLaVA-1.5 7B under greedy and nucleus decoding. Each axis represents a sub-dimension of the benchmark, and a larger area signifies better overall performance. ASCD exhibits the largest enclosed area, outperforming baseline, VCD, and ICD in most dimensions. This improvement aligns with the CHAIR and POPE findings, underscoring the benefit of selectively steering attention to reduce hallucinations.
Standard VQA Benchmarks. To verify that ASCD does not sacrifice a model’s general visual-question-answering ability, it’s evaluated on five widely-used VQA datasets. Across all three backbones and all decoding strategies, ASCD either matches or surpasses the original model on every dataset, while VCD and ICD consistently degrade performance as shown in the Appendix C.
6 Conclusion
We have shown that existing contrastive methods (e.g. VCD and ICD) inadvertently shift the internal attention distribution in MLLM, prompting us to investigate a more direct and principled way to modulate attention. We proposed an attention-steerable contrastive decoding framework that positively steers text-centric heads while negatively steering only the most critical visual tokens.
Limitations
A key limitation of ASCD is its incompatibility with FlashAttention. Because our method must dynamically modify the attention matrix at inference time, it cannot make use of the fused kernels, leading to higher memory consumption and slower decoding. A promising workaround is to distill the steering signal into the model during training: we can add an auxiliary loss – e.g., a KL-divergence term – that drives the native attention distribution to approximate the ASCD target distribution. If successful, the model would internalise the hallucination-mitigation behaviour, removing the need for on-the-fly edits and restoring FlashAttention speed-ups. We regard training-time attention regularization as a promising direction: it could distill the hallucination-robust behaviour discovered by training-free, attention-modified methods into the model itself, so that at inference the model retains this robustness while fully benefiting from FlashAttention’s speed and memory efficiency.
Appendix A Text-centric Heads Experiment Settings
The following setup applies to Section 3.3.
Finally, we aggregate these masks across all $N$ images:
$F\;=\;\sum_{i=1}^{N}M_{i},$ (8)
Appendix B Evaluation Settings
B.1 Baseline Models and Decoding Methods

### v2

@@ Introduction
Input: Image $\mathbf{I}$; text-centric heads $\mathcal{H}_{\textsc{pos}}$; critical visual-token count $\kappa_{\textsc{vis}}$; steer strengths $\alpha_{\textsc{pos}},\alpha_{\textsc{neg}}$; contrastive weight $\alpha$; truncation threshold $\beta$; MLLM with $L$ layers and $H$ heads
Output: $p_{\theta}^{\text{final}}$
1. Positive Steering Pass:
for $l\leftarrow 1$ to $L$ do
 for $h\leftarrow 1$ to $H$ do
      Compute attention matrix $\mathbf{A}^{(l)}_{h}$
    if $(l,h)\in\mathcal{H}_{\textsc{pos}}$ then
       $\mathbf{A}^{(l)}_{h}\leftarrow\mathbf{A}^{(l)}_{h}+\alpha_{\textsc{pos}}\lvert\mathbf{A}^{(l)}_{h}\rvert$
  Normalize $\mathbf{A}^{(l)}$ and continue
Obtain logits $p_{\theta}^{\text{pos}}$
[4pt]
2. Negative Steering Pass:
for $l\leftarrow 1$ to $L$ do
 for $h\leftarrow 1$ to $H$ do
      Compute attention matrix $\mathbf{A}^{(l)}_{h}$
@@ 2. Negative Steering Pass:
      Identify top-$\kappa_{\textsc{vis}}$ critical visual tokens $\mathcal{V_{\text{crit}}}$
    foreach $v\in\mathcal{V_{\text{crit}}}$ do
       $\mathbf{A}^{(l)}_{h}(v)\leftarrow\mathbf{A}^{(l)}_{h}(v)-\alpha_{\textsc{neg}}\lvert\mathbf{A}^{(l)}_{h}(v)\rvert$
  Normalize $\mathbf{A}^{(l)}$ and continue
Obtain logits $p_{\theta}^{\text{neg}}$
[4pt]
3. Contrastive Decoding and Truncation:
$p_{\theta}^{\text{raw}}\leftarrow(1+\alpha)\,p_{\theta}^{\text{pos}}-\alpha\,p_{\theta}^{\text{neg}}$
$\text{cutoff}\leftarrow\log(\beta)+\max(p_{\theta}^{\text{raw}})$
@@ 3. Contrastive Decoding and Truncation:
$p_{\theta}^{\text{final}}\leftarrow p_{\theta}^{\text{raw}}.\text{masked\_fill}(p_{\theta}^{\text{pos}}<\text{cutoff},-\infty)$
return $p_{\theta}^{\text{final}}$
Algorithm 2 Attention-Steerable Contrastive Decoding (ASCD)
Attention Steering
@@
Text-centric Head Awareness and Critical Visual Token Selection. As shown in Figure 4, we refine our method by incorporating text-centric head selection for positive steering and critical token identification for negative steering. Specifically, given the selected text-centric heads, we positively steer them by increasing their attention weights with a strength of $\alpha_{\textsc{pos}}$. Figure 5a highlights how targeted steering in text-centric heads improves the positive steering effectiveness. Simultaneously, we perform negative steering on the $\kappa_{\textsc{vis}}$ visual tokens that draw the most attention—those with the highest head-averaged score. For the current query position, we define
$s(v)=\frac{1}{H}\sum_{h=1}^{H}A^{(L)}_{h}(v),\qquad v\in\{1,\dots,V\}.$ (3)
The $\kappa_{\textsc{vis}}$ tokens with the largest $s(v)$ values form the critical set $\mathcal{V}_{\text{crit}}$; every attention entry to any $v\in\mathcal{V}_{\text{crit}}$ is then down-scaled in all heads by $\alpha_{\textsc{neg}}\lvert A^{(L)}_{h}(v)\rvert$.
Through this strategy, we deliberately obscure only the most pivotal cues – this targeted suppression is sufficient to induce a strong hallucination effect in the negative branch, leading to improved contrastive decoding compared to a blanket suppression of all visual tokens. In Figure 5b, we demonstrate the impact of selectively applying negative steering to critical visual tokens.
@@
Figure 5: Comparative effectiveness of selective attention steering. (a): Positive steering applied only to text-centric heads outperforms random or blanket head selection across various decoding strategies. (b): Negative steering focused on a small subset of critical visual tokens, integrated with contrastive decoding, significantly reduces CHAIR metrics (less hallucination) and boosts POPE scores compared to randomly suppressing visual tokens of the same number.
Integration with Contrastive Decoding with Truncation. We first compute two output distributions: $p_{\theta}^{\text{pos}}$ from the positively steered branch and $p_{\theta}^{\text{neg}}$ from the negatively steered branch. Step 3 of Algorithm 2 then fuses them through a contrastive-decoding rule with truncation, producing the final logits. This fusion amplifies visually grounded evidence while suppressing spurious text-only cues, thereby reducing hallucinations.
Experiments
Model Decoding Method CHAIRs (↓) CHAIRi (↓) POPE-Acc (↑) POPE-F1 (↑)
LLaVA-
1.5 7B
greedy Orig 53.2 13.5 85.37 84.06
VCD 56.8 15.2 84.27 83.35
ICD 52.8 13.2 83.07 80.64
PAI - - 85.82 85.79
ASCD 35.6 (33.1%) 8.6 (36.3%) 86.53 86.25
nucleus Orig 59.0 17.4 83.03 81.57
VCD 59.8 16.6 83.31 82.30
ICD 57.4 15.6 82.13 79.62
PAI - - 81.72 82.87
ASCD 43.6 (26.1%) 11.3 (35.1%) 85.75 85.07
beam Orig 54.8 15.3 85.40 84.10
VCD 58.8 16.4 84.27 83.30
ICD 52.6 13.9 83.04 80.59
PAI - - 86.33 85.89
ASCD 40.8 (25.5%) 10.1 (34.0%) 86.52 86.24
LLaVA-
NeXT 7B
greedy Orig 31.6 7.5 83.93 81.89
VCD 37.2 9.7 84.86 83.28
ICD 32.8 8.4 84.44 82.70
ASCD 21.8 (31.0%) 7.0 (6.7%) 84.85 83.40
nucleus Orig 30.4 8.0 81.74 79.61
VCD 40.4 10.4 83.55 81.95
ICD 39.4 9.9 83.67 81.98
ASCD 21.2 (30.3%) 6.7 (16.3%) 84.69 83.09
beam Orig 34.0 8.5 84.11 82.14
VCD 36.6 9.1 84.66 83.03
ICD 31.8 7.6 84.48 82.75
ASCD 21.0 (38.2%) 6.5 (23.5%) 84.91 83.48
Phi2-
SigLIP
greedy Orig 29.0 6.9 87.19 86.16
VCD 39.4 9.6 86.22 85.53
ICD 33.4 7.7 85.83 84.58
ASCD 21.8 (24.8%) 5.4 (21.7%) 87.81 86.90
nucleus Orig 36.0 9.8 85.51 84.44
VCD 36.0 8.1 85.60 84.86
ICD 37.0 9.4 84.63 83.35
ASCD 26.0 (27.8%) 8.0 (18.4%) 87.45 86.46
beam Orig 30.4 6.9 87.19 86.16
VCD 36.0 8.4 86.30 85.64
ICD 31.0 7.0 85.83 84.58
ASCD 24.6 (19.1%) 5.7 (17.4%) 87.81 86.90
LLaVA-
1.5 13B
greedy Orig 51.2 12.6 85.52 84.12
ASCD 33.0(35.5%) 8.0(36.5%) 87.78 87.40
nucleus Orig 51.4 14.5 83.95 82.54
ASCD 35.7(30.5%) 9.5(34.5%) 87.19 86.46
Qwen-
2.5-VL-
Instruct
greedy Orig 31.2 7.5 87.72 86.64
ASCD 24.4(21.8%) 6.4(14.7%) 88.91 88.27
nucleus Orig 34.8 8.6 87.34 86.14
ASCD 26.9(22.7%) 6.3(26.7%) 88.45 87.68
Table 1: CHAIR and POPE Evaluation Results. Lower CHAIRs and CHAIRi values indicate better performance in reducing hallucinations. POPE performance is reported as the mean accuracy and F1 score. The best values for each metric within a model-decoding combination are highlighted in bold. If ASCD ranks second, the best is bold while the ASCD score is underlined.
@@
To evaluate the effectiveness of our attention-steerable contrastive decoding framework in mitigating hallucinations in MLLMs, we conduct a range of experiments. This includes three diverse benchmarks—CHAIR, POPE, and MMHal-Bench—each designed to assess different aspects of object hallucinations. To ensure the broad applicability and robustness of our approach, we test it on three representative models—LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, and employ three different decoding strategies: greedy search, nucleus sampling, and beam search. Details of the experimental settings are provided in the supplementary material (Appendix, Evaluation Settings). Furthermore, we evaluate performance on standard VQA benchmarks including MMMU, MM-VET, ScienceQA, TextVQA, and GQA to verify that the proposed method preserves—rather than diminishes—the model’s original visual understanding. In addition, an extended evaluation on the larger LLaVA-1.5 13B and the modern Qwen2.5-VL-Instruct is conducted to verify scalability and architectural generality on a subset of benchmarks.
@@
It is important to note that current benchmarks for evaluating multimodal models are highly variable. For example, baseline models such as LLaVA-1.5 7B often report different metric values between different papers. Moreover, the CHAIR metric relies on random image sampling, which further complicates direct comparisons between papers. To address these issues, we faithfully reproduced both VCD and ICD using the parameters specified in their original papers and repositories, ensuring that our evaluations are conducted under consistent conditions.
POPE and CHAIR. Table 1 summarizes both caption-level (CHAIR) and VQA-style (POPE) results. Across every backbone and decoding scheme, ASCD produces the lowest CHAIR scores and the highest POPE accuracy/F1, outperforming Orig, VCD, and ICD. These gains hold for all three prompt types (random, popular, adversarial; see Appendix “Detailed POPE Results”) and remain stable when scaling from 7 B to 13 B or switching to the Qwen-VL architecture, indicating that attention steering mitigates object-level hallucinations in a model- and domain-robust manner.
MMHal-Bench. Figure 6 illustrates the radar charts of MMHal-Bench results for LLaVA-1.5 7B under greedy and nucleus decoding. Each axis represents a sub-dimension of the benchmark, and a larger area signifies better overall performance. ASCD exhibits the largest enclosed area, outperforming baseline, VCD, and ICD in most dimensions.
Standard VQA Benchmarks. To verify that ASCD does not sacrifice a model’s general visual-question-answering ability, it’s evaluated on five widely-used VQA datasets. Across all three representative backbones and all decoding strategies, ASCD either matches or surpasses the original model on every dataset, while VCD and ICD consistently degrade performance as shown in Table 2.
Additional Experiments. The supplementary material presents a runtime/memory profile and a comprehensive hyper-parameter ablation. Together, these studies confirm that ASCD adds only marginal computational overhead and remains effective across a broad range of settings, making it practical to deploy with minimal tuning effort.
Benchmark Orig VCD ICD ASCD
MM-VET (↑) 31.2 30.3 33.2 33.2
ScienceQA (↑) 67.55 67.55 67.32 69.51
GQA (↑) 61.28 59.38 59.99 61.27
TextVQA (↑) 57.82 55.07 57.66 57.91
MMMU (↑) 0.342 0.333 0.360 0.348
Table 2: Scores on five VQA-style benchmarks. The best score in each row is bold. If ASCD ranks second, the best is bold while the ASCD score is underlined.
Conclusion
We have shown that existing contrastive methods (e.g., VCD and ICD) inadvertently shift the internal attention distribution in MLLMs, prompting us to investigate a more direct and principled way to modulate attention. We propose an attention-steerable contrastive decoding framework that positively steers text-centric heads while negatively steering only the most critical visual tokens.
Appendix
Appendix A Text-centric Heads Experiment Settings
The following setup applies to "Subsection: Text-centric Heads are Model-specific".
Finally, we aggregate these masks across all $N$ images:
$F\;=\;\sum_{i=1}^{N}M_{i},$ (6)
Appendix B Extended Analysis of Text-centric Heads on Medical Data Domain
Goal.
The heads identified as text-centric were profiled on a COCO reference set containing natural photographs. This is already a very general and ideal setting for selecting text-centric heads. However, we would like to verify that this selection is not tied to the training domain but even in an extremely OOD setting, we repeat the profiling procedure on a visually dissimilar corpus of 500 NIH ChestX-ray14 images.
Protocol.
For every backbone we rerun the counting algorithm on the X-ray set, obtain a new head-frequency tensor $F_{\text{X-ray}}$, and compute the Jensen–Shannon divergence (JSD) between the COCO and X-ray distributions. For comparison we also report inter-model JSD values within each domain.
Intra-model JS Divergence (COCO vs. X-ray)
LLaVA-1.5 LLaVA-NeXT Phi-2-SigLIP
JSD 0.101 0.094 0.092
Inter-model JS Divergence on COCO
L-1.5 vs. L-NeXT L-NeXT vs. P-S L-1.5 vs. P-S
JSD 0.619 0.611 0.624
Inter-model JS Divergence on X-ray
L-1.5 vs. L-NeXT L-NeXT vs. P-S L-1.5 vs. P-S
JSD 0.604 0.572 0.620
Table 3: Jensen–Shannon divergence between text-centric-head distributions. A value below 0.1 is widely regarded as negligible drift. Abbreviations: L = LLaVA, P-S = Phi-2-SigLIP.
Result.
As shown in Table 3, all three backbones exhibit intra-model JSD $\leq 0.11$, signalling that the same set of heads remains text-centric even when the reference data shift significantly from COCO to chest X-rays. By contrast, swapping architectures while keeping the domain fixed yields JSD ${\sim}0.60$, confirming that the identity of text-centric heads is governed by model internals rather than by the reference dataset.
Appendix C Extended Evaluation: Model Scaling and Modern Architecture
Further Evaluation
We further benchmark ASCD on (i) LLaVA-1.5-13B, to test scaling effects, and (ii) Qwen2.5-VL-Instruct, one of the latest open‐source vision–language models. Table 4 reports POPE accuracy/F1 under greedy and nucleus decoding. Across both backbones and decoding schemes, ASCD yields consistent gains, demonstrating that our attention-steering approach scales beyond 3B and 7B models and transfers across different architectures.
Model Dec. Method Popular Random Adversarial Average
Acc
F1
Acc
F1
Acc
F1
Acc
F1
LLaVA-
1.5 13B
greedy Orig 0.860 0.844 0.866 0.853 0.840 0.826 0.855 0.841
ASCD 0.883 0.878 0.899 0.896 0.848 0.847 0.877 0.874
nucleus Orig 0.842 0.827 0.848 0.835 0.826 0.814 0.839 0.825
ASCD 0.876 0.867 0.887 0.880 0.851 0.845 0.871 0.864
Qwen2.5-
VL-Instruct
greedy Orig 0.878 0.865 0.886 0.876 0.868 0.857 0.877 0.866
ASCD 0.891 0.883 0.902 0.896 0.875 0.868 0.889 0.882
nucleus Orig 0.876 0.864 0.880 0.870 0.862 0.850 0.873 0.861
ASCD 0.889 0.881 0.895 0.888 0.869 0.860 0.884 0.876
Table 4: POPE benchmark results on a larger (LLaVA-1.5 13B) and a novel (Qwen2.5-VL-Instruct) backbone. Bold numbers mark improvements of ASCD over the corresponding original model under the same decoding strategy; gray rows highlight ASCD.
Visualization of Text-centric Heads Distribution of LLaVA-1.5 13B and Qwen2.5-VL-Instruct
Figure 7 visualizes the head-frequency heatmap for LLaVA-1.5 13B and Qwen2.5-VL-Instruct. Both models reproduce the sparse, high-contrast pattern already observed for the previous backbones: a small set of heads appear in dark colors, indicating that they are marked text-centric in the vast majority of reference samples, while the remainder receive little or no votes.
Figure 7: Text-centric-head frequency maps. Each pixel denotes the probability that a given layer–head is selected as text-centric when profiled on 500 COCO images. Darker colours highlight a handful of heads that dominate textual attention, revealing the same long-tailed, head-level anisotropy seen in previous models.
Appendix D Detailed POPE Results
Table 5 presents the full POPE breakdown by different type—random, popular, and adversarial—for each backbone and decoding scheme. We report accuracy and F1. Across all settings, ASCD consistently matches or surpasses the baselines, with the largest gains typically appearing on the random split.
Model Dec. Method Popular Random Adversarial Average
Acc
F1
Acc
F1
Acc
F1
Acc
F1
@@ Appendix D Detailed POPE Results
Appendix E Evaluation Settings
Baseline Models and Decoding Methods